OpenStudy (anonymous):

What is the simplified form of the expression 3[14 / (4^2 - 3^2) - 5] A) -9 B) 27 C) 21 D) 6

OpenStudy (anonymous):

@bruhhh

OpenStudy (displayerror):

Is your function \[3 \times \left( \frac{14}{4^2 - 3^2 -5} \right)\] or \[3 \times \left( \frac{14}{4^2 - 3^2} -5 \right)\] ?

OpenStudy (anonymous):

thats actually a division sign, not fraction line lol

OpenStudy (displayerror):

It's the same thing. \[\frac{4}{2} = 4 \div 2 = 2\]

OpenStudy (anonymous):

I think it'd be the 2nd one

OpenStudy (displayerror):

Depending on how your problem is formatted (either the first or second one), you would get two different answers. If the problem was given as you had typed it, then I would think it's the second one as well.